How many pints of a 9% cleaning solution must be mixed with 4 pints of a 15% cleaning solution to give a 12% solution?
Answer by stanbon(75887)

How many pints of a 9% cleaning solution must be mixed with 4 pints of a 15% cleaning solution to give a 12% solution?
-----------
Equations:
active + active = active
0.09x + 0.15*4 = 0.12(x+4)
---
Multiply thru by 100 to get:
9x + 15*4 = 12x + 12*4
----
-3x = -3*4
x = 4 pints (amt of 9% solution needed in the mix)
---
Cheers,
Stan H.